7u+3r=724 and 3u+7r=1276
How much is each? (r and u)

Answer by Alan3354(69443)

7u+3r=724
3u+7r=1276
--------
21u + 9r = 2172 eqn 1 times 3
21u +49r = 8932 eqn 2 times 7
---------------- Subtract
-40r = -6760
r = 169
---------
7u + 3*169 = 724
u = 31